OpenStudy (anonymous):

5(x+y)=20+3x what does y =

OpenStudy (campbell_st):

y = (20 + 3x)/5 - x y = 4 -2x/5

OpenStudy (kymber):

Distribute, Subtract 5x, Divide by 5, Y = -2/5x + 4